Database last updated on Mon, 02 Feb 2026 03:40:49 +0000 by localhost

Critical warnings from b-files' comments for days between 2024024 and 2024024 (0 sec, 3 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
hg:Lamp not detected f1= 0 f4= 24251B02424.026TORON#02605:38:071-1
hg:Lamp not on... hg test terminated F1= 0 F4= 24251B02424.026TORON#02605:38:071-1
hg:HG failed31B02424.026TORON#02605:39:57failed1-1

Instrument constant changes for the week before 2024024 (1 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 16 rows)

Non-critical warnings from b-files' comments for days between 2024024 and 2024024 (0 sec, 3 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
dp:Doing dispersion with internal HG.299B02424.026TORON#02611:48:261-1
rs:HOME key pressed199B02424.026TORON#02618:44:031-1
pd:Going to menu199B02424.026TORON#02618:44:091-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhgrssluv
024TORONTO#026sch026,Waiting until ZA =,Waiting until ZA = 170 now 123.3,Waiting until ZA = 170 now 125.1,Waiting until ZA = 170 now 126.9,Waiting until ZA = 170 now 128.7,Waiting until ZA = 170 now 130.5,Waiting until ZA = 170 now 132.2,Waiting until ZA = 170 now 134,Waiting until ZA = 170 now 135.7,Waiting until ZA = 170 now 137.4,Waiting until ZA = 170 now 139,Waiting until ZA = 170 now 140.7,Waiting until ZA = 170 now 142.3,Waiting until ZA = 170 now 143.8,Waiting until ZA = 170 now 145.3,Waiting until ZA = 170 now 146.7,Waiting until ZA = 170 now 148.1,Waiting until ZA = 170 now 149.4,Waiting until ZA = 170 now 150.6,Waiting until ZA = 170 now 151.6,Waiting until ZA = 170 now 152.6,Waiting until ZA = 170 now 153.4,Waiting until ZA = 170 now 154.1,Waiting until ZA = 170 now 154.7,Waiting until ZA = 170 now 155,Waiting until ZA = 170 now 155.22826626291
(1 sec)

Dead time (0 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B02424.02628#0263128